OpenStudy (anonymous):

How can you represent the terms of a sequence explicitly? How can you represent them recursively? Provide mathematical examples to support your opinions.

OpenStudy (anonymous):

a sequence can be written by a summation

OpenStudy (anonymous):

Explicitly means you define it as a function of \(n\). \[ a_n = f(n) \]E.G.\[ a_n = 54 - 5n \] Recursively means you define it in terms of other parts of the sequence E.G. \[ a_n = a_{n-1} + a_{n+1} \cdot a_{n/2} \]

OpenStudy (anonymous):

or simply calculated in other words you can have 1+2+3+4 or \[\sum_{n=0}^\infty n+1\]